OpenStudy (anonymous):

Math12: Geometric Sequences The first three terms of a geometric sequence are sqrt[4]{3},sqrt[8]{3}, and 1. What is the fourth term?

OpenStudy (anonymous):

I would really appreciate if you guys can show me steps by steps

OpenStudy (mimi_x3):

Can you please right it clearly. Is it: \[\sqrt{4*3}?\]

OpenStudy (mimi_x3):

write* lol

OpenStudy (anonymous):

.\[\sqrt[4]{3} , \sqrt[8]{3}, and 1.\]

OpenStudy (anonymous):

sorry about that :p

OpenStudy (mimi_x3):

Well, find the common ratio first.. \[\large \frac{T_{2}}{T_{1}} = \frac{\sqrt[8]{3}}{\sqrt[4]{3}} \] \[\large \frac{T_{3}}{T_{2}} = \frac{1}{\sqrt[8]{3}} \]

OpenStudy (mimi_x3):

Then find the general term; the formula is: \[\large T_n = ar^{n-1}\] To find the fourth term; sub in n = 4; where a is the first term of the sequence and r is the common ratio.

OpenStudy (anonymous):

cool! thanks you!! that was helpful :)